OpenStudy (anonymous):

On the graph of the equation 2x + 5y = –30, what is the value of the y-intercept?

jimthompson5910 (jim_thompson5910):

plug in x = 0 and solve for y to get the y-intercept

OpenStudy (anonymous):

You are wanting to find the value where this equation crosses the y axis. When does it cross the y axis? It crosses the y axis when x = 0. So, substitute 0 for x in the equation and then solve.

OpenStudy (anonymous):

so the y-intercept would be 6?

OpenStudy (anonymous):

2(0) + 5y = -30 0 + 5y = -30 5y = -30 y = -30/5 y = -6 you were close, but forgot the negative

OpenStudy (anonymous):

or 2x + 5y = –30, put that in y=mx+b form, so 5y=-2x-30, you need to get y by itself so divide everything by 5, -30/5=-6 , so -6, you Are just dividing the b value by the y value.